OpenStudy (anonymous):

Determine the number of possible triangles, ABC, that can be formed given angle A = 30°, a = 4, and b = 6. a. 0 b. 1 c. 2

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

can you help

OpenStudy (amoodarya):

one way to solve is : using law of cosine find "c" then you can find number of triangle \[b^2+c^2-a^2=2(b)(c) \cos(A)\\36-16+c^2=2(6)(c)\cos(30)\]

OpenStudy (amoodarya):

or using law of sine \[\frac{ \sin(A) }{ a }=\frac{ \sin(B) }{ b}\\\frac{\sin30}{4}=\frac{\sin(B)}{6}\\sin(B)=\frac{3}{4}\]

OpenStudy (anonymous):

how can i get my answer from 3/4?

OpenStudy (amoodarya):

B=?

OpenStudy (anonymous):

6

OpenStudy (amoodarya):

can you find B?

OpenStudy (amoodarya):

angle of B ?

OpenStudy (anonymous):

i dont know how

OpenStudy (amoodarya):



OpenStudy (amoodarya):

B=48.5 or B=180-48.5

OpenStudy (anonymous):

131.5?

OpenStudy (anonymous):

thats not one of my answers?

OpenStudy (anonymous):

is the answer 1?

OpenStudy (anonymous):

i think its 2

OpenStudy (amoodarya):

then c=? a=30 b=131.5 , b=48.5

OpenStudy (amoodarya):

number of possibilities is ?